OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1070! If three times the number of one-step equations minus twice the number of two-step equations is equal to 110, how many one-step equations auditioned today?

OpenStudy (anonymous):

x = one step equations y = two step equations x + y = 1070 -->y = 1070 - x 3x - 2y = 110 now sub 1070 - x in for y in the 2nd equation, and solve for x (your 2 step equations)